A rectangular pool has dimensions of 40 ft. and 60 ft. The pool has a patio area around it that is the same width on all sides.If the patio area equals the area of the pool, how wide is the distance from the pool edge to the patio edge?
If x represents the width of the patio, then which of the following equations could be used to solve the word problem?
Answer by checkley77(12844) 
 40*60=2,400 ft^2 is the area of the pool.
(40+2x)(60+2x)=2*2,400
2400+120x+80x+4x^2=4,800
4x^2+200x+2,400-4,800=0
4x^2+200x-2,400=0
4(x^2+50x-600)=0
4(x-10)(x+60)=0
x-10=0
x=10 ans. for the width of the patio.
Proof:
(40+2*10)(60+2*10)=2*2,400
(40+20)(60+20)=4,800
60*80=4,800
4,800=4,800
| |
|